OpenStudy (anonymous):

Given that x, 4, x+6 are consecutive terms of a geometric series, find: (a) the possible values of x So I got x=2 but part from trying every number to see I don't know of a method to find the possible values. I don't know how many values they want so if there is a method that shows all the possible values that would be better.

OpenStudy (welshfella):

Geometric series so common ratio = second term / first = third / second so 4 / x = (x + 6) / 4 solve for x

OpenStudy (welshfella):

that rearranges to the quadrataic equation x^2 + 6x - 16 = 0

OpenStudy (welshfella):

can you solve that?

OpenStudy (welshfella):

x = 2 is one of the possible values

OpenStudy (welshfella):

you need to factor x^2 + 6x - 16 you want 2 numbers whose product is -16 and whose sum = + 6

OpenStudy (welshfella):

have you learned how to solve these type of equations?

OpenStudy (anonymous):

Yes I have done quadratics before. I will try that now quickly

OpenStudy (anonymous):

I got the right thing! Thank you very much!

OpenStudy (welshfella):

yw